### 23 Jan 2024 v1

# UniHDA: Towards Universal Hybrid Domain Adaptation of Image Generators

Abstract
Generative domain adaptation has achieved remarkable progress, enabling us to adapt a pre-trained generator to a new target domain. However, existing methods simply adapt the generator to a single target domain and are limited to a single modality, either text-driven or image-driven. Moreover, they are prone to overfitting domain-specific attributes, which inevitably compromises cross-domain consistency. In this paper, we propose UniHDA, a unified and versatile framework for generative hybrid domain adaptation with multi-modal references from multiple domains. We use CLIP encoder to project multi-modal references into a unified embedding space and then linear interpolate the direction vectors from multiple target domains to achieve hybrid domain adaptation. To ensure the cross-domain consistency, we propose a novel cross-domain spatial structure (CSS) loss that maintains detailed spatial structure information between source and target generator. Experiments show that the adapted generator can synthesise realistic images with various attribute compositions. Additionally, our framework is versatile to multiple generators, e.g., StyleGAN2 and Diffusion Models.
Figure 1: Application of UniHDA. Given a pre-trained source generator and multiple target domains (either in text or image modalities), UniHDA adapts the generator to a hybrid target domain that blends all characteristics and maintains strong consistency with source domain.
1 Introduction
Figure 2: Linear interpolation between multi-modal direction vectors. We represent the domain shift by the direction vector from source embedding to the target (e.g., Crying or Happy). Linear interpolation of them during training will result in a smooth traversal.
Benefiting from tremendous success of modern image generator [8, 1, 31, 25], generative domain adaptation has achieved remarkable progress during the past few years. Typically, it aims to adapt a pre-trained generator to a new target domain while preserving the variation in the source domain, e.g., from the human domain to the baby domain. Depending on the modality of references, generative domain adaptation can be categorized into two schools: text-driven [5, 17] and image-driven [15, 14, 19, 40, 35, 16].
Despite their promising results, there are still some limitations. A major limitation of existing methods is that they only support adaptation from a source domain to individual target domain. These methods fail to directly adapt a human generator to more practical real-world scenarios like angry person with the style of sketch or a combination of Johnson and Elena (Fig. 1). For more general purposes, previous methods also fail with multi-modal references, e.g., smiling Joker given the non-smiling Joker image and the smile text.
Another limitation of existing approaches is that, in the pursuit of faithful adaptation, they are prone to overfit domain-specific attributes, especially in image-driven scenarios [19]. This inevitably compromises cross-domain consistency and impedes the inheritance of the diversity from the source generator.
In this paper, we propose UniHDA, a Universal framework for generative Hybrid Domain Adaptation with multi-modal references from multiple domains. UniHDA facilitates the reference of individual image and text prompt simultaneously and blends the attributes from diverse target domains to create a hybrid domain. To enable multiple modalities, we leverage pre-trained CLIP [24] to project multi-modal references into a unified embedding space and represent the domain shift by the direction vector from the source embedding to the target embeddings.
To achieve hybrid domain adaptation, we draw inspiration from the compositional capabilities in the latent space of StyleGAN [6, 27, 36]. We demonstrate that a semantically meaningful linear interpolation between two direction vectors in CLIP’s embedding space can uncover favorable compositional capabilities, as shown in Fig. 2. In light of this intriguing finding, we linearly interpolate direction vectors of multiple target domains to obtain the direction vector corresponding to the hybrid domain that semantically integrates attributes from all target domains.
Furthermore, we introduce a novel cross-domain spatial structure loss (CSS) to preserve the consistency between source and target generator by maintaining detailed spatial structure information. Concretely, we leverage pre-trained DINO-ViT [4, 21] to encode generated images into patch tokens with fine-grained spatial information. For cross-domain consistency, we maintain the correspondence between source and target tokens with contrastive learning [20].
We conduct experiments for a wide range of source and target domains to validate the effectiveness of our method. Results demonstrate that the adapted generator can synthesise realistic images with various attribute compositions. In addition, we show that UniHDA is agnostic to the type of generators (e.g., StyleGAN2 [8, 9, 10] and Diffusion models [7, 11]). Our contributions are as follows:
  • We propose a unified and versatile framework for generative domain adaptation, which enables multi-modal references and hybrid target domain, e.g., text-text, image-image, and image-text.
  • We demonstrate strong compositional capabilities of direction vectors in CLIP’s embedding space. Taking advantage of it, we propose to linearly interpolate the direction vectors for hybrid domain adaptation.
  • We propose a cross-domain spatial structure loss to maintain fine-grained spatial structure information between source and target generator, which significantly improves cross-domain consistency.
  • We demonstrate successful multi-modal hybrid domain adaptation to diverse new domains for various generators and illustrate its advantage over other methods.
2 Related Work
Text-driven Generative Domain Adaptation. Text-driven domain adaptation involves using a textual prompt to shift the domain of a pre-trained model toward a new domain. For example, Style-NADA [5] presents a local direction CLIP [24] loss to align the embeddings of the generated images and text. Based on Style-NADA, Domain Expansion (DE) [18] proposes to expand the generator to jointly model multiple domains with texts.
Image-driven Generative Domain Adaptation. Image-driven generative domain adaptation refers to the adaptation of a pre-trained image generator to a new target domain using a limited number of training images. Due to the scarcity of training images, prior researches  [15, 14, 19, 40, 35, 16, 33, 39, 41] often integrate additional regularization terms to prevent overfitting. For instance, CDC [19] introduces the instance distance consistency loss to maintain the distance between different instances in the source domain. Although these researches have made significant strides in generative domain adaptation, they heavily rely on the discriminator, making it challenging to handle hybrid domain adaptation.
Generative Hybrid Domain Adaptation. To combine multiple domains and achieve hybrid domain adaptation, several domain adaptation methods propose to train a separate generative model per domain and combine their effects in test-time, e.g., Style-NADA [5]. Although having potential for hybrid domain adaptation, it doubles the model size and training time due to the requirement for training multiple models separately. Domain Expansion (DE)  [18] proposes to expand the generator to jointly model multiple domains via decompose latent space. However, it requires the source dataset (e.g., FFHQ [8]) for regularization, which significantly increases training time. Furthermore, it relies on the semantic latent space of the generator (e.g., StyleGAN [8] and DiffAE [23]) for hybrid domain adaptation, limiting its applicability to a broader range of generators. Recently, FHDA [13] proposes few-shot hybrid domain adaptation and introduces a directional subspace loss to achieve it. Differently, we focus on multi-modal references with the text and 1-shot image, which offers greater flexibility and versatility.
Disentanglement in Generative Models.  As observed in StyleGAN [8], the latent space is essentially a linear subspace. Recent works [6, 27, 36, 28, 34, 22, 32, 30] propose to find individual latent factors for image variations. Among them, SeFa [27] computes the eigenvalues of the transformation matrix to find the latent directions. As for diffusion models, DiffAE [23] explores the possibility of using DPMs for representation learning and seeks to extract a meaningful and decodable representation of an input image via autoencoding. Drawing inspiration from the disentanglement in the latent space of them, we demonstrate that a semantically meaningful linear interpolation between the direction vectors in CLIP’s embedding space can similarly uncover favorable compositional capabilities, which facilitates us to achieve hybrid domain adaptation.
3 Method
Figure 3: Overview of UniHDA with multi-modal direction loss $\mathcal{L}_{direct}$ and cross-domain spatial structure loss $\mathcal{L}_{\text{CSS}}$. Utilizing CLIP image encoder and text encoder, $\mathcal{L}_{direct}$ encourages $G_{\mathcal{T}}$ to faithfully acquire domain-specific characteristics with multi-modal references. To facilitate diversity inherited from $G_{\mathcal{S}}$, $\mathcal{L}_{\text{CSS}}$ improves cross-domain consistency by maintaining detailed spatial structure information. The red solid line represents positive pairs, while red dashed lines represent negative pairs.
3.1 Multi-Modal Hybrid Domain Adaptation
We start with a pre-trained generator $G_{\mathcal{S}}$ (e.g., StyleGAN [8, 9, 10] and Diffusion model [7, 29])), that maps from noise $z$ to images in a source domain $\mathcal{S}$. Given a new target domain $\mathcal{T}$ referenced by texts [5, 38, 12, 11] or images [15, 14, 19, 40, 35, 16], generative domain adaptation aims to adapt $G_{\mathcal{S}}$ to yield a target generator $G_{\mathcal{T}}$, which can generate images similar to domain $\mathcal{T}$.
Despite the promising results of existing methods, a major limitation of them is that they only support adaptation from the source domain to individual target domains and fail to directly adapt the generator to the hybrid domain which blends the characteristics of multiple domains. Furthermore, they fail with multi-modal adaptation driven by texts and images simultaneously.
For more general purposes, we propose multi-modal hybrid domain adaptation. Given $N$ domains $\{\mathcal{T}_{i}\}_{i=1}^{N}$ with 1-shot image $\{Y_{i}\}$ and $M$ domains $\{\mathcal{T}_{j}\}_{j=1}^{M}$ with the text prompt $\{P_{j}\}$, it aims to adapt the source generator $G_{\mathcal{S}}$ to $G_{\mathcal{T}}$ that models the hybrid domain $\mathcal{T}=\{\mathcal{T}_{i}\}\cup\{\mathcal{T}_{j}\}$ and generates images with integrated characteristics. To the end, we introduce UniHDA, a unified and versatile framework for multi-modal hybrid domain adaptation, as shown in Fig. 3.
3.2 Multi-modal Direction Loss
To enable multiple modalities, we leverage pre-trained CLIP [24] to encode text-image references into a unified semantic embedding space. Drawing inspiration from CLIP-based methods [5, 38, 12, 11], we represent the domain shift as the direction vector $\Delta f_{dom}$ from the source embedding and the target embedding. For image reference $Y_{i}$ and its CLIP embedding $f_{i}$, the domain shift is calculated by
$\begin{gathered}\Delta f_{dom}=f_{i}-\overline{f_{s}},\end{gathered}$ (1)
where $\overline{f_{s}}$ is the mean embedding of several samples generated by $G_{\mathcal{S}}$. For text prompt $P_{j}$ and its CLIP embedding $f_{j}$,
$\begin{gathered}\Delta f_{dom}=f_{j}-\widetilde{f_{s}},\end{gathered}$ (2)
where $\widetilde{f_{s}}$ is the embedding of the source text prompt.
To adapt $G_{\mathcal{S}}$, we initialize a new generator $G_{\mathcal{T}}$ from $G_{\mathcal{S}}$ and finetune it by aligning the sample-shift direction $\Delta f_{samp}$ with the domain-shift direction $\Delta f_{dom}$. Formally,
$\begin{gathered}\Delta f_{samp}=f_{t}-f_{s},\\ \mathcal{L}_{direct}=1-\frac{\Delta f_{samp}\cdot\Delta f_{dom}}{\left\|\Delta f_{samp}\right\|\left\|\Delta f_{dom}\right\|},\end{gathered}$ (3)
where $f_{s}$ and $f_{t}$ are the embeddings of samples generated by $G_{\mathcal{S}}$ and $G_{\mathcal{T}}$ with the same noise.
3.3 Linear Composition of Direction Vectors
To achieve the hybrid domain adaptation, we draw inspiration from the compositional capabilities in the latent space of StyleGAN [6, 27, 36]. We illustrate that a linear interpolation between two direction vectors in the embedding space of CLIP, which is semantically meaningful, reveals promising compositional capabilities. As shown in Fig. 2, we can smoothly interpolate between two direction vectors calculated by distinct target prompts and source prompt “photo”, resulting in a gradual adaptation toward the target domain.
In light of this intriguing finding, we employ linear interpolation on the direction vectors of multi-modal target domains, to derive a consolidated direction vector representing the hybrid domain that seamlessly integrates all attributes semantically. For given domain coefficients $\{\alpha_{i}\}$ and $\{\alpha_{j}\}$, we obtain the direction vector
$\begin{gathered}\Delta f_{dom}=\sum_{i=1}^{N}\alpha_{i}(f_{i}-\overline{f_{s}})+\sum_{j=1}^{M}\alpha_{j}(f_{j}-\widetilde{f_{s}}),\end{gathered}$ (4)
which represents the domain shift between the hybrid domain and source domain. We then substitute Eq. 4 into Eq. 3 to adapt $G_{\mathcal{S}}$ to the hybrid domain.
3.4 Cross-domain Spatial Structure Loss
Albeit the direction loss achieves multi-modal hybrid domain adaptation, the adapted generator are prone to overfit domain-specific attributes. This exacerbates when it comes to image-image and image-text scenarios owing to the scarcity of the images. To address this issue, we introduce a novel cross-domain spatial structure loss (CSS) to enhance cross-domain consistency, ensuring the preservation of intricate spatial structural information between the source and target generator.
Specifically, we leverage the pre-trained DINO-ViT [4, 21] to encode the generated images into patch tokens, containing detailed spatial structural information. DINO-ViT is self-supervised trained to focus on the distinction between subjects of the same class [26], which facilitates us to maintain the cross-domain consistency. Motivated by contrastive learning [20], we reduce the distance between the positive token pairs at the same position and push away the negative token pairs at the other positions by
$\begin{gathered}\mathcal{L}_{\text{CSS}}=-\sum_{i}\log\frac{\exp(v_{i}^{t}\cdot v_{i}^{s})}{\sum_{j}\exp(v_{i}^{t}\cdot v_{j}^{s})},\end{gathered}$ (5)
where $v_{i}^{t}$ and $v_{j}^{s}$ are the $i$-th and $j$-th tokens in the last layer of DINO-ViT from $G_{\mathcal{T}}$ and $G_{\mathcal{S}}$ respectively. The dot mark $\cdot$ represents cosine similarity.
Overall, our training loss consists of two terms, i.e., the multi-modal direction loss $\mathcal{L}_{direct}$ to achieve multi-modal hybrid domain adaptation and the cross-domain spatial structure loss $\mathcal{L}_{\text{CSS}}$ to maintain cross-domain consistency:
$\begin{gathered}\mathcal{L}_{overall}=\mathcal{L}_{direct}+\lambda\mathcal{L}_{\text{CSS}},\end{gathered}$ (6)
where we use $\lambda=5$ in our experiments.
Figure 4: Image-image hybrid domain adaptation. We compare the results of FHDA [13], NADA [5] and UniHDA (Ours) with the same noise. FHDA and NADA generate images with poor cross-domain consistency, leading to a limited diversity. In contrast, UniHDA alleviates overfitting and maintains strong cross-domain consistency.
Table 1: Quantitative results for image-image domain adaptation.
Method Taylor-Elena Hulk-Wooden Johnson-Comic CS-I ($\uparrow$) SCS ($\uparrow$) CS-I ($\uparrow$) SCS ($\uparrow$) CS-I ($\uparrow$) SCS ($\uparrow$) FHDA 0.685 0.576 0.635 0.659 0.640 0.679 NADA 0.684 0.579 0.624 0.575 0.647 0.642 Ours 0.699 0.738 0.649 0.707 0.656 0.764
Method Baby-Red hair Baby-Happy Big-Blue eyes
CS-T ($\uparrow$) SCS ($\uparrow$) CS-T ($\uparrow$) SCS ($\uparrow$) CS-T ($\uparrow$) SCS ($\uparrow$)
DE 0.163 0.638 0.160 0.580 0.195 0.662
NADA 0.179 0.661 0.170 0.642 0.186 0.731
Ours 0.186 0.744 0.175 0.757 0.197 0.765
Table 1: Quantitative results for image-image domain adaptation.
Table 2: Quantitative results for text-text domain adaptation.
4 Experiments
4.1 Experimental Setting
Methodology. We demonstrate the versatility of UniHDA on multi-modal hybrid domain adaptation, i.e., text-text, image-image, and image-text. To show the generator-agnostic nature of UniHDA, we apply it to two well-known generators, i.e., StyleGAN2 [9] and Diffusion model [11]). Following previous generative domain adaptation literatures  [5, 38, 15, 14, 18, 19, 40, 35, 16], we use StyleGAN2 for comparisons in most experiments.
Datasets. We conduct the experiments for a wide range of source and target domains to validate the effectiveness of UniHDA. Following previous work, we consider FFHQ [8], AFHQ Dog [3], and LSUN Church [37] as the source domains. We adapt the pre-trained models to diverse hybrid domains driven by the text prompt and 1-shot image, where we set the domain coefficients in Eq. 4 as 0.5. Unless stated otherwise, we operate on 1024 × 1024 resolution images for both source and target domains.
Evaluation Metrics. One important aspect to evaluate generative domain adaptation is the preservation of domain-specific characteristics. Following  [26], we use CLIP Score (CS-T and CS-I) for text-text and image-image adaptation respectively. Concretely, CS-T is measured by the average cosine similarity between prompt and generated images’ embedding. CS-I is the average pairwise cosine similarity between CLIP embeddings of generated and real images. Here we use average CS-T or CS-I of multiple domains. Another important evaluation is the cross-domain consistency of the source domain. To measure it, we adopt Structural Consistency Score (SCS) [35] to evaluate the spatial structural consistency between source and target generator.
Figure 5: Text-text hybrid domain adaptation. We compare the results of DE [18], NADA [5] and UniHDA (Ours) with the same noise. UniHDA exhibits desirable performance to acquire characteristics from hybrid target domain and maintain robust cross-domain consistency.
4.2 Image-image Hybrid Domain Adaptation
Fig. 4 shows the qualitative results of baselines and UniHDA for image-image adaptation. As shown in the figure, FHDA [13] suffers from severe model collapse and generates images with a limited diversity due to the scarcity of image references. While NADA [5] mitigates overfitting to a certain extent, its cross-domain consistency remains poor, resulting in the generation of similar images. In contrast, UniHDA maintains strong consistency and effectively generates images with characteristics of the hybrid domain.
We also quantitatively compare UniHDA with baselines. As shown in Tab. 2, ours clearly outperforms other methods. We achieve better CS-I and SCS, indicating that generated images effectively integrate domain-specific attributes and preserve primary characteristics of the source domain.
4.3 Text-text Hybrid Domain Adaptation
Fig. 5 shows the qualitative results of the baselines and UniHDA for text-text adaptation, all of which start from the same source domain FFHQ [8] to the combinations of individual domains. Since the adaptation is conducted solely along one projection direction of the latent code, Domain Expansion (DE) [18], to some extent, does not fully capture the characteristics of the target domain, e.g., baby (Row 1 and Row 2). Furthermore, DE does not maintain robust cross-domain consistency, e.g., the chin of the person in the upper-right corner and background artifacts in the Row 3. This is primarily due to the characteristics bias introduced when projecting the latent code to the base subspace, making the images appear less authentic. The main problem of NADA [5] is overfitting. Hard-to-learn characteristics, e.g., baby (Row 1 and Row 2) and big eyes (Row 3) may be overshadowed by other overfitted ones. In contrast, UniHDA (Ours) exhibits desirable performance to generate images with integrated characteristics while maintaining robust consistency with the source domain.
Similar to Sec. 4.2, we also compare UniHDA with the baselines quantitatively. As shown in Tab. 2, ours clearly outperforms the baselines, which are consistent with qualitative results in Fig. 5. For CS-T, UniHDA significantly outperforms other methods, indicating that generated images effectively integrate multiple characteristics from distinct domains. Furthermore, UniHDA achieves better SCS, which effectively maintains cross-domain consistency compared with baseline methods.
Method Extra Data Dependency Model Dependency 2-domain 10-domain
size ($\downarrow$) time ($\downarrow$) size ($\downarrow$) time ($\downarrow$)
DE Source dataset $*{}^{*}$ 24M 20h 24M 20h
NADA - Multi-Models 48M 4min 240M 20min
Ours - Single-Model 24M 2min 24M 2min
  • *** indicates that DE needs a semantic latent space of the generator.
Table 3: Comparison with DE [18] and NADA [5]. For extra data dependency, DE needs source dataset (e.g., FFHQ [8]) that significantly increases the training time. Besides, DE requires a semantic latent space, which limits the generalization for more generators. For hybrid $N$-domains, NADA utilizes $N$ adapted models and necessitates $N$ times the model size and training time.
Figure 6: Image-text hybrid domain adaptation. Compared with NADA [5], UniHDA maintains strong cross-domain consistency.
Figure 7: Results of UniHDA for image-text hybrid domain adaptation on dogs [3], and churches [37]. The adapted generator can generate visually pleasant images with robust cross-domain consistency.
Figure 8: Results of UniHDA with DiffusionCLIP [11]. We replace the training objective of DiffusionCLIP with our proposed $\mathcal{L}_{direct}$ and $\mathcal{L}_{\text{CSS}}$. Results demonstrate that UniHDA is agnostic to the type of generators, allowing for its broader application on the diffusion model.
Figure 9: Ablation for our proposed $\mathcal{L}_{\text{CSS}}$ on single domains, which significantly alleviates overfitting and improves cross-domain consistency.
4.4 Image-text Hybrid Domain Adaptation
Fig. 6 shows the results of image-text adaptation. As depicted in Sec. 4.2, NADA is susceptible to overfitting, which retains poor cross-domain consistency. In contrast, UniHDA (Ours) achieves robust cross-domain consistency and generates diverse images in all scenarios. Furthermore, we showcase the results of image-text hybrid domain adaptation on dogs [3], and churches [37] in Fig. 7. We can clearly observe that it achieve desirable results across various target domains. More results are included in Appendix.
4.5 Comparison with Existing Methods
In  Sec. 4.2 and  Sec. 4.3 , we find that while DE [18] and NADA [5] compromise cross-domain consistency, UniHDA (Ours) retains strong consistency and successfully captures the attributes of target domains. Besides, UniHDA also surpasses them in terms of efficiency, e.g., model size and training time as shown in Tab. 3. NADA trains a separate generative model per domain and interpolates their parameters in test-time, which necessitates multiple times the model size and training time. Although DE avoids cross-model interpolation, it heavily relies on the large source dataset (e.g., FFHQ [8]) for regularization during training process, resulting in a significant increase in training time. In contrast, UniHDA circumvents these issues, which enables the completion of the adaptation within single generator and only two minutes.
Additionally, DE relies on the semantic latent space of the generator (e.g., StyleGAN [8] and DiffAE [23]) for hybrid domain adaptation, limiting its applicability to a broader range of generators. Conversely, UniHDA is not constrained by the type of generators, allowing for its broader application across various generators.
4.6 Generalization on Diffusion Models
In this section, we demonstrate that UniHDA is agnostic to the type of generative models and can easily generalize to diffusion models. Specifically, we apply UniHDA on DiffusionCLIP [11] and replace the training objective of DiffusionCLIP with our proposed $\mathcal{L}_{direct}$ and $\mathcal{L}_{\text{CSS}}$. As shown in Fig. 8, the results work as well as those on StyleGAN2, which integrate the characteristics from multiple target domains and maintain robust consistency with the source domain. This verifies desirable generalization ability and versatility of UniHDA.
4.7 Effective of CSS Loss
We conduct the ablation study to evaluate the effects of our proposed Cross-domain Spatial Structure (CSS) loss on single domains. As shown in Fig. 9, the results without $\mathcal{L}_{\text{CSS}}$ suffer from overfitting and have very limited cross-domain consistency, e.g., distorted backgrounds in Row 1 and 3. Benefited from CSS, the generated images maintain consistency with the source images in terms of spatial structure, thereby inheriting the diversity from the source domain. More results are included in Appendix.
5 Conclusion & Limitation
In this paper, we propose UniHDA, a unified and versatile framework for multi-modal hybrid domain adaptation. To enable multiple modalities, we leverage CLIP encoder to project multi-modal references into a unified embedding space. For hybrid domain, we demonstrate the compositional capabilities of direction vectors in CLIP’s embedding space and linearly interpolate direction vectors of multiple target domains. In addition, we propose a new cross-domain spatial structure loss to improve cross-domain consistency by preserving intricate spatial information between the source and target generators. We believe that our work is an important step towards generative domain adaptation since we have demonstrated that the source generator can be effectively adapted to a hybrid domain with multi-modal references.
While UniHDA effectively realizes multi-modal hybrid domain adaptation, it also has the limitation. To encode both image and text into a shared embedding space, we utilize pre-trained CLIP during training time, which might bring potential bias for some domains. Nevertheless, we believe that the exploration of the novel task are significant for future work and solutions could be integrated into UniHDA to eliminate the bias.
Supplementary Material
Figure 10: More qualitative results to verify the effectiveness of our proposed $\mathcal{L}_{\text{CSS}}$.
Figure 11: Hybrid domain adaptation in 3D generator. To show the versatility of UniHDA, we apply it on the popular 3D-aware generator, EG3D [2].
Figure 12: Hybrid domain adaptation from AFHQ cat to incompatible domains, i.e. lion, rabbit, and tiger.
Figure 13: Hybrid domain adaptation from AFHQ dog to incompatible domains, i.e. lion, rabbit, and tiger.
Figure 14: Effect of different pre-trained image encoders for CSS on image-image hybrid domain adaptation.
Figure 15: Effect of different pre-trained image encoders for CSS on text-text hybrid domain adaptation.
Figure 16: Effect of different pre-trained image encoders for CSS on image-text hybrid domain adaptation.
Figure 17: More results of UniHDA with DiffusionCLIP.
Figure 18: More results of image-image hybrid domain adaptation. The source image is in the top-left corner, and the first row and column consist of training images.
Figure 19: More results of text-text hybrid domain adaptation. The source image is in the top-left corner, and the first row and column consist of text prompts.
Figure 20: More results of image-text hybrid domain adaptation. The source image is in the top-left corner. The first row and column consist of training images and text prompts respectively.
Figure 21: More results of image-text hybrid domain adaptation on AFHQ dog.
6 Appendix
$\mathcal{L}_{\text{CSS}}$ Elena Dog Church
0.569 0.429 0.617
0.748 0.671 0.752
Table 4: Quantitative study of our $\mathcal{L}_{\text{CSS}}$, which significantly improves cross-domain consistency and achieves better SCS score.
6.1 More Ablation of CSS Loss
As depicted in Sec. 4.7, our proposed $\mathcal{L}_{\text{CSS}}$ significantly alleviates overfitting and improves cross-domain consistency. We conduct the quantitative study in Tab. 4, which aligns with the earlier qualitative results in Fig. 9. Results with $\mathcal{L}_{\text{CSS}}$ achieve better SCS score, indicating that they maintain stronger consistency with the source domain. Additionally, we show more qualitative results in Fig. 10 to verify the effectiveness of UniHDA.
6.2 Hybrid Domain Adaptation of 3D Generator
To verify the versatility of UniHDA, We apply it on the popular 3D-aware image generation method, EG3D [2]. Specifically, we replace the discrimination loss with our proposed $\mathcal{L}_{direct}$ and $\mathcal{L}_{\text{CSS}}$ for hybrid domain adaptation. As shown in Fig. 11, the results effectively integrate the attributes and preserve the characters and poses of source domain.
6.3 Results of Incompatible Domain Adaptation
To verify the generalization ability of UniHDA, we additionally conduct experiments for hybrid domain adaptation on incompatible domains, i.e. from cat to rabbit. As shown in Fig. 12 and Fig. 13, we start from AFHQ cat and dog to incompatible domains respectively. We can observe that the results effectively integrate the attributes of the corresponding domain and maintain robust consistency with source domain.
6.4 Effect of Encoder for CSS
We conduct experiments on pre-trained ViT, MViTv2, and Dinov2 to explore the impact of different image encoders for CSS. As shown in Fig. 14, Fig. 15, and Fig. 16, we showcase the results of image-image, text-text, and image-text respectively. We can observe that all of them improve the consistency with source domain compared with the baseline approach. Furthermore, they exhibit a similar qualitative style, which demonstrates that our CSS is agnostic to different pre-trained image encoders.
6.5 More Results for DiffusionCLIP
To demonstrate the versatility of UniHDA, we apply it on DiffusionCLIP in Sec. 4.6. As shown in Fig. 17, we showcase more results for DiffusionCLIP including image-image, text-text, and image-text. All results achieve hybrid domain adaptation and preserve strong corss-domain consistency.
6.6 More Qualitative Results
We apply UniHDA to adapt the generator on FFHQ to more hybrid domains, i.e., text-text, image-image, and image-text. As shown in Fig. 18, Fig. 19, and Fig. 20, UniHDA successfully generates images with integrated characteristics from multiple target domains and maintains robust consistency with the source domain. Besides, we showcase more results of hybrid domain adaptation from AFHQ dog in Fig. 21.
6.7 Implement Details
Following the setting of previous generative domain adaptation methods [5, 18], we utilize the batch size of 4 and ADAM Optimizer with a learning rate of 0.002 for all experiments during training. A training session typically requires 300 iterations in 2 minutes, which significantly reduces training time compared with adversarial methods for generative domain adaptation. Note that we conduct all experiments on a single NVIDIA RTX 4090 GPU. The code will be open source.
For experiments on FFHQ, we generate images with 1024 × 1024 resolution. As for AFHQ-Dog and LSUN-Church, we operate on 512 × 512 and 256 × 256 resolution images respectively.